# The Holomorphic Extension Property for $k$-du Bois Singularities

### v2

@@ 3.4. An Extension Criterion for Differentials on a Resolution of Singularities
    (3.10) $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0}))\leq n-j-p-2$
@@ 4.1. Work of Kovács-Schwede-Smith
Note that $\mathrm{Supp}~{}E_{2}^{p,q}\neq 0$ if $j=-\dim\Sigma,...,0$, whence $\dim\mathrm{Supp}~{}E_{2}^{p,q}=0$ if $p<\dim Z-q$ by Theorem 2.1(iv). Thus $E_{\infty}^{p,q}=0$ in this range, which gives the desired vanishing on $X$. ∎
Proof of Theorem 1.1. By assumption, $\underline{\Omega}_{X}^{0}$ has no higher cohomology, and $\dim\mathrm{Supp}~{}\mathscr{O}_{X}<n-2$ since $X$ is normal. Therefore logarithmic extension holds in degree $n$ by Lemma 4.1 and Proposition 1.5. By [KS21, Theorem 1.5], logarithmic extension holds in all degrees $0\leq p\leq n$. ∎
@@ 5.1. Consequences of Logarithmic Extension
$\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))\leq-(j+p+2)$
@@
$\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))=0,$
@@
Let $X$ be a normal variety of dimension $n$ with at worst du Bois singularities, and let $\mathrm{IC}_{X}=(\mathcal{I}C_{X},F_{\bullet},IC_{X})$ be the intersection Hodge module. Then
@@
Let $X\subset Y$ be a (locally) closed embedding into a smooth complex manifold $Y$. Let $N_{0}=(\mathcal{N}_{0},F_{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$ be the logarithmic mixed Hodge module. By construction, $W_{n}\mathbb{D}_{Y}(\mathcal{N}_{0})=\mathbb{D}_{Y}(W_{n}\mathcal{N}_{0})$. Since $\mathrm{gr}_{-1}^{F}\mathrm{DR}(-)$ is an exact functor, there is an exact sequence
$\begin{split}\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))&\to\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}(\mathrm{gr}_{n}^{W}\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))\\ &\to\mathscr{H}^{1}\mathrm{gr}_{-1}^{F}\mathrm{DR}(W_{n-1}\mathbb{D}(\mathcal{N}_{0})(-n)).\end{split}$
The last term must be zero for degree reasons, and so by Corollary 5.1 we get the additonal vanishing $\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathrm{gr}_{n}^{W}\mathbb{D}_{Y}(\mathcal{N}_{0})(-n))=0$. By (3.3), we have an isomorphism $\mathrm{gr}_{n}^{W}\mathbb{D}_{Y}(\mathcal{N}_{0})(-n)\cong\mathrm{gr}_{n}^{W}\mathcal{N}_{0}$. The claim follows from the isomorphism
@@ 5.3. Log Forms to Holomorphic Forms
Proof of Theorem 1.7. Let $\mathrm{IC}_{X}$ and $N_{0}$ be the intersection and logarithmic Hodge modules of the normal variety $X$. We note there is a canonical morphism of (mixed) Hodge modules $M_{X}\to N_{0}$ obtained by the isomorphism $W_{n}N_{0}\cong\mathrm{IC}_{X}$ [KS21, Proof of (9.8.1)]. The problem is local, so let $X\subset Y$ be a closed embedding of $X$ into a ball $Y\cong\mathbb{C}^{n+c}$, and let $H\subset Y$ be a generic hyperplane section. For each $p$, there is a commutative diagram of exact sequences
[…]
@@
$\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})=\mathscr{H}^{-\dim\Sigma}\mathrm{gr}_{-((n-1)-\dim\Sigma)}\mathrm{DR}(\mathcal{IC}_{X\cap H}).$
For dimension reasons, this term vanishes. Therefore the morphism
$\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\pi_{*}\Omega_{\widetilde{X}}^{p}\to\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$
is an isomorphism by the 4-lemma. The theorem follows since $H$ was chosen to be general. ∎
@@
6. Holomorphic Extension for (Weakly) $k$-du Bois Singularities
6.1. Rational v.s. Weakly $k$-du Bois Singularities
@@ 6. Holomorphic Extension for (Weakly) $k$-du Bois Singularities
Therefore, it seems interesting to ask how closely (weakly) $k$-du Bois singularities are from having rational singularities. Even for $k>0$, there is a disconnect. The following affine cone examples were described in [Tig23], see also [SVV23, §7.5].
@@
Here is an example of an affine cone singularity which is rational but is not $k$-du Bois for some $11\leq k\leq n-$ 1. The example is essentially given in [BTLM97, §4.1] and is related to the failure of Bott vanishing for non-projective spaces. Let $Y\subset\mathbb{P}^{4}$ be a smooth quadric hypersurface. Let $L=\mathscr{O}_{Y}(1)$. Then the affine cone $X$ has rational singularities since $Y$ is a Fano variety. The cone is also weakly 1-du Bois. To see this [^3] Since $Y$ is a hypersurface, then so will the affine cone $X$. Therefore this follows from [MOPW21], we consider the cohomology sequence of
@@
In summary: weakly $k$-du Bois singularities need not have rational singularities, nor are rational singularities weakly $k$-du Bois for all $k$. What is interesting about Example 6.1 is that the 3-fold singularity is weakly $3$-du Bois, and yet holomorphic extension does not hold in degree 3. This is because $\mathscr{H}^{0}\underline{\Omega}_{X}^{2}$ is also not reflexive in this case, a necessary condition for holomorphic extension to hold for $k$-du Bois singularities.
@@ 6.2. Isolated Singularities and Depth
To further highlight what Example 6.1 tells us, we consider the relationship between $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$, $\pi_{*}\Omega_{\widetilde{X}}^{p}$, and $\mathrm{depth}(\mathscr{O}_{X})$. Let $X$ be a normal variety with isolated singularities. If $X$ is weakly $k$-du Bois, then Theorem 2.1(vii) implies a short exact sequence
@@
6.3. A criterion for Holomorphic Extension for Weakly $k$-du Bois Singularities
@@
Suppose $X$ is a normal complex algebraic variety of dimension $n$. If $X$ is weakly $k$-du Bois and $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ is reflexive, then holomorphic extension holds in degree $k+1$.
@@
By Proposition 1.3 and the quasi-isomorphism $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}\cong_{\mathrm{qis}}\underline{\Omega}_{X}^{k}$, the assumption on $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ implies $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathbb{D}_{X}(\underline{\Omega}_{X}^{k})\leq-(j+2)$ for $j\geq-n+1$. By looking exact triangle of Theorem 2.1(ix), dualizing, and taking cohomology, we get the vanishing $R^{n-1}\pi_{*}\Omega_{\widetilde{X}}^{n-k}(\log E)=0$. By the residue exact sequence, $R^{n-1}\pi_{*}\Omega_{E_{i}}^{n-k-1}=0$ for each smooth component $E_{i}$ of $E$. Since the problem is local, we can shrink $X$ and assume that $H^{n-1}(E_{i},\Omega_{E_{i}}^{n-k-1})=0$. The $E_{i}$ are smooth and projective (see §2.1), so $H^{0}(E_{i},\Omega_{E_{i}}^{k})=0$, too. This implies the inclusion $H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{k+1})\hookrightarrow H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{k+1}(\log E))$ is an isomorphism. Since $X$ is du Bois, the claim follows from Theorem 1.2. ∎
@@
Let $X$ be a normal complex variety with weakly $k$-du Bois singularities for $k\geq\mathrm{codim}_{X}(\Sigma)-1$, where $\Sigma$ is the singular locus of $X$. If $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for each $0\leq p\leq k$, then
@@
For isolated singularities, the assumption on $k$ implies $\pi_{*}\omega_{\widetilde{X}}$ is reflexive by Theorem 6.3. Therefore the claim (6.1) holds by Proposition 1.3 and Grauert-Rimenschneider vanishing. More generally, let $X\subset Y$ be a locally closed embedding into a smooth manifold $Y$ and let $H$ be a general hyperplane section of $Y$. By Theorem 2.1(vi), $X\cap H$ has weakly $k$-du Bois singularities, and it is clear that $\mathscr{H}^{0}\underline{\Omega}_{X\cap H}^{p}$ is reflexive whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive. Therefore induction and the isomorphism

### v3

@@ 3.4. An Extension Criterion for Differentials on a Resolution of Singularities
    (3.10) $\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{N}_{0}))\leq n-j-p-2$
@@ 4.1. Work of Kovács-Schwede-Smith
Note that $\mathrm{Supp}~{}E_{2}^{p,q}\neq 0$ if $j=-\dim\Sigma,...,0$, whence $\dim\mathrm{Supp}~{}E_{2}^{p,q}=0$ if $p<\dim\Sigma-q$ by Theorem 2.1(iv). Thus $E_{\infty}^{p,q}=0$ in this range, which gives the desired vanishing on $X$. ∎
Proof of Theorem 1.1. By assumption, $\underline{\Omega}_{X}^{0}$ has no higher cohomology, and $\dim\Sigma\leq n-2$ since $X$ is normal. Therefore logarithmic extension holds in degree $n$ by Lemma 4.1 and Proposition 1.4. By [KS21, Theorem 1.5], logarithmic extension holds in all degrees $0\leq p\leq n$. ∎
@@ 5.1. Consequences of Logarithmic Extension
$\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{p}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))\leq-(j+p+2)$
@@
$\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))=0,$
@@
Let $X$ be a normal variety of dimension $n$ with at worst Du Bois singularities, and let $\mathrm{IC}_{X}=(\mathcal{I}C_{X},F^{\bullet},IC_{X})$ be the intersection Hodge module. Then
@@
Let $X\subset Y$ be a (locally) closed embedding into a smooth complex manifold $Y$. Let $N_{0}=(\mathcal{N}_{0},F^{\bullet},W_{\bullet},\mathrm{rat}~{}N_{0})$ be the logarithmic mixed Hodge module. By construction, $\mathrm{gr}_{-n}\mathbf{D}_{Y}(\mathcal{N}_{0})=\mathbf{D}_{Y}(W_{n}\mathcal{N}_{0})$. Since $\mathrm{gr}_{-1}^{F}\mathrm{DR}(-)$ is an exact functor, there is an exact sequence
$\begin{split}\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))&\to\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}(\mathrm{gr}_{n}^{W}\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))\\ &\to\mathscr{H}^{1}\mathrm{gr}_{-1}^{F}\mathrm{DR}(W_{n-1}\mathbb{D}(\mathcal{N}_{0})(-n)).\end{split}$
The last term must be zero for degree reasons, and so by Corollary 5.1 we get the additonal vanishing $\mathscr{H}^{0}\mathrm{gr}_{-1}^{F}\mathrm{DR}(\mathrm{gr}_{n}^{W}\mathbf{D}_{Y}(\mathcal{N}_{0})(-n))=0$. By (3.3), we have an isomorphism $\mathrm{gr}_{n}^{W}\mathbf{D}_{Y}(\mathcal{N}_{0})(-n)\cong\mathrm{gr}_{n}^{W}\mathcal{N}_{0}$. The claim follows from the isomorphism
@@ 5.3. Log Forms to Holomorphic Forms
Proof of Theorem 1.7. Let $\mathrm{IC}_{X}$ and $N_{0}$ be the intersection and logarithmic Hodge modules of the normal variety $X$. We note there is a canonical morphism of (mixed) Hodge modules $\mathcal{IC}_{X}\to N_{0}$ obtained by the isomorphism $W_{n}N_{0}\cong\mathrm{IC}_{X}$ [KS21, Proof of (9.8.1)]. The problem is local, so let $X\subset Y$ be a closed embedding of $X$ into a ball $Y\cong\mathbb{C}^{n+c}$, and let $H\subset Y$ be a generic hyperplane section. For each $p$, there is a commutative diagram of exact sequences
00{0}00{0}${N_{H|Y}^{*}\otimes_{\mathscr{O}_{H}}\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})}$${N_{H|Y}^{*}\otimes_{\mathscr{O}_{H}}\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{N}_{0}|_{H})}$${\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})}$${\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\mathscr{H}^{-(n-p)}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{N}_{0})}$${\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})}$${\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{N}_{0}|_{H})}$${\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})}$${\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{N}_{0}|_{H})}$
@@
$\mathscr{H}^{-(n-p)+1}\mathrm{gr}_{-p+1}^{F}\mathrm{DR}(\mathcal{IC}_{X\cap H})=\mathscr{H}^{-\dim\Sigma}\mathrm{gr}^{F}_{-((n-1)-\dim\Sigma)}\mathrm{DR}(\mathcal{IC}_{X\cap H}).$
For dimension reasons, this term vanishes, and we see the map
$\mathscr{O}_{H}\otimes_{\mathscr{O}_{Y}}\pi_{*}\Omega_{\widetilde{X}}^{p}\to(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p}$
is surjective.
Next, we follow an argument in [GK14a]. Let $E=\sum E_{i}$ be the exceptional divisor of $\pi$, and let $\alpha$ be section in $H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{p}(\log E))$. By definition, there are integers $m_{i}\in\left\{0,1\right\}$ such that $\gamma$ is a section in $H^{0}(\widetilde{X},\Omega_{\widetilde{X}}^{p}(\sum m_{i}E_{i}))$. Consider the diagram
[…]
We assume again the assumption is true in dimension $\dim X-1$. In particular, the image $\beta(\alpha(\gamma))$ factors through $H^{0}(\widetilde{H},\Omega_{\widetilde{H}}^{p})$. Supposing by contradiction that the $m_{i}$ are not all 0, this implies the existence of a nonzero section of $H^{0}(\widetilde{H},N_{\widetilde{H}|\widetilde{X}}^{*}\otimes\Omega_{\widetilde{X}}^{p-1}\otimes(\sum m_{i}E_{i}))$. Since the pullback of $N_{H|X}$ is just $N_{\widetilde{H}|\widetilde{X}}$, this implies $H^{0}(\widetilde{H},\Omega_{\widetilde{H}}^{p-1}\otimes\mathscr{O}(\sum m_{i}E_{i}))$. By induction and our choice of $p$, it is clear the morphism
$(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p-1}\to(\pi|_{H})_{*}(\Omega_{\widetilde{H}}^{p-1}\otimes\mathscr{O}(\sum m_{i}E_{i}))$
is surjective. We have also seen the morphism $\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p-1}\to(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p-1}$ is surjective for our choice of $p$. This produces a non-zero section of $H^{0}(\widetilde{H},\Omega_{\widetilde{H}}^{p-1}(\sum m_{i}E_{i})|_{\widetilde{H}})$ which vanishes under $\beta$. We can continue to iterate this process until we receive a non-zero section of $H^{0}(\widetilde{H},\mathscr{O}_{\widetilde{H}}(\sum m_{i}E_{i}))$. In fact, we can find for each $E_{j}$ with $m_{j}=1$ that
$H^{0}(\widetilde{H},\mathscr{O}(m_{i}E_{i})|_{E_{j}})\neq 0.$
This is a contradiction to the negativity lemma [GKK10, Proposition 7.5]. ∎
@@
6. Holomorphic Extension for (Weakly) $m$-Du Bois Singularities
6.1. Rational v.s. Weakly $m$-Du Bois Singularities
@@ 6. Holomorphic Extension for (Weakly) $m$-Du Bois Singularities
Therefore, it seems interesting to ask how closely (weakly) $m$-Du Bois singularities are from having rational singularities. Even for $k>0$, there is a disconnect. The following affine cone examples were described in [Tig23], see also [SVV23, §7.5].
@@
Here is an example of an affine cone singularity which is rational but is not $m$-Du Bois for some $11\leq k\leq n-$ 1. The example is essentially given in [BTLM97, §4.1] and is related to the failure of Bott vanishing for non-projective spaces. Let $Y\subset\mathbb{P}^{4}$ be a smooth quadric hypersurface. Let $L=\mathscr{O}_{Y}(1)$. Then the affine cone $X$ has rational singularities since $Y$ is a Fano variety. The cone is also weakly 1-Du Bois. To see this [^3] Since $Y$ is a hypersurface, then so will the affine cone $X$. Therefore this follows from [MOPW21], we consider the cohomology sequence of
@@
In summary: weakly $m$-Du Bois singularities need not have rational singularities, nor are rational singularities weakly $m$-Du Bois for all $k$. What is interesting about Example 6.1 is that the 3-fold singularity is weakly $3$-Du Bois, and yet holomorphic extension does not hold in degree 3. This is because $\mathscr{H}^{0}\underline{\Omega}_{X}^{2}$ is also not reflexive in this case, a necessary condition for holomorphic extension to hold for $m$-Du Bois singularities.
@@ 6.2. Isolated Singularities and Depth
To further highlight what Example 6.1 tells us, we consider the relationship between $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$, $\pi_{*}\Omega_{\widetilde{X}}^{p}$, and $\mathrm{depth}(\mathscr{O}_{X})$. Let $X$ be a normal variety with isolated singularities. If $X$ is weakly $m$-Du Bois, then Theorem 2.1(vii) implies a short exact sequence
@@
6.3. A criterion for Holomorphic Extension for Weakly $m$-Du Bois Singularities
To summarize, the weaker version of the higher Du Bois property does not appear to detect the holomorphic extension property. However, we can remedy this by requiring $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ to be reflexive. The following is an extension of Theorem 1.2 for weakly $m$-Du Bois singularities.
@@ 6.3. A criterion for Holomorphic Extension for Weakly $m$-Du Bois Singularities
Let $X$ be a normal complex algebraic variety of dimension $n$ with singular locus $\Sigma$. Suppose $X$ is weakly $m$-Du Bois, where $k=\mathrm{codim}_{X}(\Sigma)-1$. If $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ is reflexive, then $\pi_{*}\Omega_{\widetilde{X}}^{k+1}$ is reflexive.
@@
For isolated singularities, we are checking $\pi_{*}\omega_{\widetilde{X}}$ is reflexive when $X$ is weakly $(n-1)$-Du Bois and $\mathscr{H}^{0}\underline{\Omega}_{X}^{n-1}$ is reflexive. By assumption, the quasi-isomorphism
$\mathscr{H}^{0}\underline{\Omega}_{X}^{n-1}\cong\mathbf{R}\pi_{*}\Omega_{\widetilde{X}}^{n-1}(\log E)(-E)$
implies $\dim\mathrm{Supp}~{}R^{j}\pi_{*}\Omega_{\widetilde{X}}^{1}(\log E)\leq n-j-2$. In particular, $R^{n-1}\pi_{*}\Omega_{\widetilde{X}}^{1}(\log E)=0$. This gives the additional vanishing $R^{n-1}\pi_{*}\Omega_{E_{(1)}}^{1}=0$ by the residue exact sequence. We may pass to cohomology by shrinking $X$ as necessary; by Hodge theory, this gives $\pi_{*}\Omega_{E_{(1)}}^{n-1}=0$. This implies $\pi_{*}\omega_{\widetilde{X}}\hookrightarrow\pi_{*}\omega_{\widetilde{X}}(E)$ is an isomorphism. Since $X$ is assumed to be Du Bois, this proves the result by Theorem 1.1.
In order to prove $\pi_{*}\Omega_{\widetilde{X}}^{\mathrm{codim}_{X}(\Sigma)}$ is reflexive, it is sufficient to prove the stronger claim
$\dim\mathrm{Supp}~{}\mathscr{H}^{j}\mathrm{gr}_{-\dim\Sigma}^{F}\mathrm{DR}(\mathcal{IC}_{X})\leq-(j+2).$
For $k=n-1$ (i.e., isolated singularities), this is equivalent to $\pi_{*}\omega_{\widetilde{X}}$ being reflexive. We can therefore proceed by induction as in the proof of Theorem 1.7. Specifically, the higher Du Bois property is preserved by general hyperplane [SVV23, Theorem A], and the reflexivity of $H^{0}\underline{\Omega}_{X}^{\mathrm{codim}_{X}(\Sigma)}$ is also preserved by hyperplane. As a result, we get the desired support condition except possibly when $j=-1$; but this is only relevant when $X$ has isolated singularities. ∎
@@
Let $X$ be a normal complex variety with weakly $m$-Du Bois singularities for $k\geq\mathrm{codim}_{X}(\Sigma)-1$, where $\Sigma$ is the singular locus of $X$. If $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for each $0\leq p\leq k$, then
@@
For isolated singularities, the assumption on $k$ implies $\pi_{*}\omega_{\widetilde{X}}$ is reflexive by Theorem 6.3. Therefore the claim (6.1) holds by Proposition 1.3 and Grauert-Riemenschneider vanishing. More generally, let $X\subset Y$ be a locally closed embedding into a smooth manifold $Y$ and let $H$ be a general hyperplane section of $Y$. By Theorem 2.1(vi), $X\cap H$ has weakly $m$-Du Bois singularities, and it is clear that $\mathscr{H}^{0}\underline{\Omega}_{X\cap H}^{p}$ is reflexive whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive. Therefore induction and the isomorphism